0. Approximate a root of f using (a) a fisx nplo 1 Consider the function f)r method. and (b) Newtons Method

numerical analysis Newton and fixed point iteration method
Answer #1

a) We keep applying the function cos(z) o(z) COS T for the initial point = 0.5 to get the below sequence of iterations:

n x_n
0 0.5
1 0.877583
2 0.639012
3 0.802685
4 0.694778
5 0.768196
6 0.719165
7 0.752356
8 0.730081
9 0.74512
10 0.735006

Thus, an approximate fixed point (i.e. solution to cos(x) =x COS T) is 0,735

b) Using f(x)=\cos(x)-x\Rightarrow f'(x)=-\sin(x)-1

We get the iteration scheme: cos(x) - 2r ㅡ sin (2)-1 rn+1 = o(%) where o(x) = x-f(r) f(x) Tn +1-o(m) where ψ(z

That is, cos(x) - ir o(r) = r + sin(z) +1

Which means r sin() r +cos(x) - r TTCOST sin(x) 1 sin(r) +

Which is 0(1)-!sinir) tools(r sin(r) + 1

Applying this iteration scheme, we get:

n x_n
0 0.5
1 0.755222
2 0.739142
3 0.739085
4 0.739085
5 0.739085

So that the root is 0.739085